CMS Technical Incident Management

 Home | Contact Us | Search

  Technical Coordination




What's New
Service Request
Support Forum

The purpose of this Web site is to  provide a repertoire of technical incidents with some explanation on how they have been tackled.  Resources are also provided here to help report and resolve problems and suggest improvements.

Main Categories

  • Power Cuts

       What to do if there is a power cut

  • Network Interruption
  • Cooling Interruption
  • DSS Actions

        temp sms list for information

        Responsibles and Delay Table

  • Gas News
We are flushing about 400 l/h Ar-CO2 (40-60) through the racks of the CSC + End-cap (5 distribution units). The chambers are bypassed. Two volume buffers are connected to the rack in ME+1, to be able regulate pressures better.
  • Miscellaneous
  • Lift Issues
  • Access Issues

Detailed information

Power Cut Logs

Network Logs

Cooling Logs

Gas Problem Logs

Miscellaneous Logs





  • Power Cut Action Manual (in French; need help for translation) (.pdf)
  • Network Troubleshooting Manual (.doc)
  • Cooling Actions Manual (.doc)
  • Gas Troubleshooting Manual (.doc)
  • Miscellaneous Actions Manual (.doc)

Current Issues


Cooling un-operational in SX5 from 16:35; temp high, and we still do not have a T sensor DSS cut off; DCS works nevertheless; though still to understand why some rack temps are not displayed. short report


Both Lifts un-operational last evening, but ok now. TCR - 72201 was called yesterday; lifts were fixed; but panels were left there. That created a confusion this morning. The shift assumed that the lifts were not working due to the signs on the panels, and three hours were lost.

Cooling problems as well - no sms messages received - to be checked on Monday 23.7

Look for the temp excursion sample from S1D09

Highlights of the week

EB SM Installation - 11 modules inserted

YB0 Services Installation -

Magnet Cryo test begin

Important Phone Numbers:


16 0408

TIM 16 5000
GLIMOS 16 0202
SLIMOS 16 5500
UX Czar 16 2541
US Czar 16 4076
TCR 72201
Gas Piquet 160954


Scheduled Tests and Stoppages

1.A cold test will take place on July 30th for AGI insulation: No Access Restriction

UXC55: When will the final ventilation system be operational in UXC55 - At the end of July BUT the Plug needs to be closed


P5: Diesel groups will be changed for more powerful ones by the end of August. Two week activity during which the power backup will be supplied by the CERN network. The switch will have to be done manually. Possible start date: end September. An IS37 to be issued


DSS tests in USC55 between DSS system and turbine racks will take place next week. DSS tests UXC55 racks should start once smoke detection in racks is operational

UXC55: Ramses test with radioactive source will start on 26/07 D Perrin

UXC55: water has to be pumped from the jack holes: Fire brigade
13.Drainage gallery: Ineo has to install lighting

Request from TS-IC-HM to maintain crane PR784 during W38-39 (17-28/09):

USC55: TS-IC-HM to define a clearance area on top of Cryogenics system

Monday 23.7  
Tuesday 24.7  
Wednesday 25.7  
Thursday 26.7  
Friday 27.7  
Saturday 28.7  
Other Magnet He tests

LHC 4-5 Sectors  closed for ~ 3- 4 months






Last modified: 07/15/07